# FUSAR-KLIP: Towards Multimodal Foundation Models for Remote Sensing

### v3

@@ 1 Introduction
In recent years, foundational models built around the Transformer have established a dominant position in the field of computer vision, demonstrating outstanding general representation capabilities through self-supervised paradigms such as masked image modeling or contrastive learning[10876763],[10490262]. Inspired by this, a number of foundational models targeting the characteristics of Earth observation have rapidly emerged in the field of remote sensing. To address the need for high-precision sensing, RingMo[diao2025ringmo] and SpectralGPT[10490262] have achieved deep adaptation to small targets and multispectral features through customized masked modeling and 3D spectral reconstruction. To address the challenges of wide-area generation and data sparsity, MetaEarth[yu2024metaearth] and AlphaEarth[brown2025alphaearth] have introduced generative diffusion and spatiotemporal embedding field techniques, successfully extending the model’s capabilities from passive interpretation to active simulation and multi-source assimilation, forming a diversified development pattern centered around the core needs of Earth observation[10607962],[zhou2024vlgfm],[10770756],[soni2025earthdial],[liuRandomFeatures].
While the aforementioned work has achieved record-breaking performance metrics on specific tasks, its theoretical foundation still faces significant challenges. As Muttenthaler et al. pointed out, a fundamental misalignment exists between intelligent model representations and human cognition: existing deep neural networks, although excelling in local feature matching, often fail to capture hierarchical knowledge structures like humans[muttenthaler2025aligning]. This means that current models lack the deep cognitive abilities that align with human logic, resulting in significantly limited robustness and generalization when faced with out-of-distribution data.
@@
This cognitive disparity has further evolved into a dual barrier of modal heterogeneity and mechanism intricacy in the field of synthetic aperture radar (SAR)[10965804]. Although SAR possesses irreplaceable all-weather strategic value, its imaging, based on microwave coherent mechanisms, is heavily affected by speckle noise and geometric distortion[10638186]. When existing foundation models are migrated to the SAR domain, they struggle to understand the geometric structure and physical properties of ground objects through non-intuitive scattering phenomena, leading to a sharp decline in generalization performance[zhou2025fifty]. Therefore, constructing a multimodal foundation model that adapts to SAR imaging characteristics and possesses geographic cognitive capabilities has become a critical scientific imperative in the current field of remote sensing intelligence.
@@
1) Geospatial Priors Remain Underutilized: Geographic information, a core element in Earth sciences, is crucial for remote sensing. However, most current SAR interpretation studies remain at the visual level, neglecting geographic attributes. This stems from the extension of natural image paradigms—focused mainly on texture and structure[Liu2016MRELBP]—as well as the closed nature of SAR data sources, with many studies still relying on outdated datasets like MSTAR from the 1990s[AFRL_MSTAR]. As shown in Table I, our review of 16 mainstream SAR datasets across tasks such as detection, segmentation, and recognition reveals a widespread omission of geographic metadata. This absence hinders regional analysis, terrain understanding, and spatial reasoning, ultimately limiting progress toward cognitive-level SAR interpretation.
2) Knowledge Bottleneck: Existing remote sensing VLM studies face a significant knowledge bottleneck in text construction. On the one hand, approaches that convert detection, recognition, or segmentation annotations into templated text [10504785],[wei2025sarlang] generate descriptions that are semantically sparse and structurally repetitive, covering only shallow attributes (e.g., category and location) while neglecting complex elements such as landform, regional function, or spatial structure. This low-information text causes different images to collapse into similar semantic spaces, weakening alignment quality and fine-grained discrimination, as shown in Fig. 2(a). On the other hand, automatic text acquisition strategies, which are effective in natural image or optical remote sensing domains (e.g., web crawling, video subtitles, or AI-generated captions), fail in the SAR domain due to speckle noise and the need for expert knowledge. The resulting AI-generated text is often semantically vague or distorted Fig. 2(b). Together, these issues highlight that SAR multimodal foundational model suffers from a dual bottleneck: text derived from manual annotations lacks semantic depth, and reliable large-scale automatic acquisition remains infeasible, severely restricting scalability in open-world applications.
@@
FUSAR-MAP[shi2021object] 2021 45 C Single 1 1,024 610 TIF Segmentation
SSDD[XU2025440] 2021 633 C,X Multi $\sim1$5 500 1,160 JPG Detection
@@
FUSAR-SHIP[hou2020fusar] 2022 271 C Single 1 512 16,144 TIFF Detection
@@
SAMPLE[lewis2019sar] 2019 14 X Single 0.3 128 2,732 PNG Classification
@@
PolarSAR-Seg[wang2022air] 2022 36 C Multi 8 512 500 TIFF Segmentation
@@
In terms of text data acquisition, this paper introduces a hierarchical cognitive chain-of-thought (HCoT) instruction to simulate the human interpretation process of SAR imagery. HCoT guides GPT-4.1 [OpenAIChatGPT] to progressively incorporate multi-dimensional knowledge—such as geographic context, regional priors, SAR imaging principles, and target scale perception—enabling semantic reasoning from global to local, and from general to domain-specific knowledge. Furthermore, the constructed multi-scale image semantics driving mechanism guides the model to generate semantic expressions that connect the upper and lower scales on large, medium, and small scale images, and establish cognitive coherence. Under HCoT guidance, the language model significantly enhances its SAR interpretation ability, generating more informative and coherent textual descriptions to support multimodal representation learning.
In terms of multimodal alignment modeling, this paper proposes FUSAR-KLIP, a dual-tower multimodal framework employing Vision Transformer (ViT)[dosovitskiy2021an] and BERT[devlin2019bert] as visual and language encoders. The model jointly optimizes image-text contrastive loss (ITC), image-text matching loss (ITM), and masked language modeling loss (MLM) to construct a unified cross-modal embedding space bridging low-level perception and high-level semantics. To mitigate semantic deviations in HCoT-guided generated text, we introduce a self-consistent iterative optimization (SCIO) module that enhances alignment accuracy and stability through a closed-loop self-supervised strategy of screening, filtering, and reconstruction.
@@ 2.1 Paradigm Evolution of Remote Sensing Interpretation
In the supervised learning stage, large-scale manual annotations enabled significant progress in recognition, detection, and segmentation tasks, but the prohibitive cost and limited semantic coverage of annotations severely constrained generalization across diverse scenes[shi2021object].
Unimodal self-supervised learning (SSL) alleviated this bottleneck by mining latent structures from unlabeled data, producing robust representations with minimal fine-tuning and laying the foundation for domain-specific models[10713915]. The RingMo series pioneered the introduction of mask learning into remote sensing, optimizing feature extraction of dense small targets in remote sensing images through targeted masking strategies[diao2025ringmo]. RingMo-Aerial further introduced a frequency-enhanced attention mechanism and affine transformation contrastive learning, solving the challenges of tilted viewpoints and multi-scale occlusion in aerial remote sensing. Addressing the 3D characteristics of multispectral data, SpectralGPT innovatively employs 3D masking and multi-target reconstruction strategies, overcoming the limitation of processing only RGB data and achieving effective capture of spectral sequence information[10490262]. In the generative direction, MetaEarth, through a self-cascaded framework and noise sampling strategy, achieved global-scale, unbounded resolution remote sensing image generation, driving the leap from perception to simulation[yu2024metaearth]. RS-vHeat takes a different approach, using a heat conduction physics model to simulate feature diffusion, significantly improving computational efficiency while maintaining the global receptive field[hu2025rs]. Furthermore, AlphaEarth proposed the concept of Embedding Fields, addressing the label sparsity problem by assimilating multi-source spatiotemporal data[brown2025alphaearth]. Nevertheless, unimodal SSL remains largely restricted to low-level perceptual cues such as texture and geometry, offering limited capacity for higher-order semantic reasoning[10915556],[10955443].
@@ 2.2 Research on Unimodal Self-supervised Foundational Models in SAR Images
SARATR-X uses a two-step SSL method with multi-scale gradient features to establish a high-performance SAR image target recognition foundational model[10856784]. Yang et al. proposed the SARDet-CL, combining feature enhancement with a SSL method constrained by imaging mechanisms, and achieved advanced performance in downstream detection tasks[10994283]. Li et al. proposed SAR-JEPA, which constructed a self-supervised pre-training foundational model for SAR target recognition tasks from the perspective of overcoming speckle noise[LI2024326]. Ren et al. introduced multi-image factor SSL to promote directional feature learning and obtain generalized features, enhancing the performance in terrain classification tasks[10496083]. Wang et al. constructed a SAR multi-task foundation model based on cross-domain continuous pre-training[10726860]. Pei et al. proposed a two-stage SAR image pre-training method based on SSL to improve the accuracy of target classification[10269665]. MSFA proposed a multi-stage filter augmentation pre-training framework for large-scale RGB and SAR data, which performed well when transferred to detection tasks[li2024sardet100k].
@@ 2.3 Research on Multimodal VLMs in Remote Sensing Field
Existing work primarily focuses on optical imagery. For instance, RemoteCLIP is based on the CLIP framework by converting class labels in public remote sensing datasets into templated text, establishing a foundational model that achieves leading performance across many typical tasks[10504785]. SkyEyeGPT developed a multimodal remote sensing command dataset and designed a two-stage tuning strategy to enhance conversational capabilities[zhan2025skyeyegpt]. VHM generated the HqDC dataset based on the large language model Gemini, easing the issue of model hallucination[pang2025vhm]. SkyScript utilized structured geographic information from OpenStreetMap to generate semantic text for optical images, improving the quality of cross-modal alignment[wang2024skyscript]. GeoChat expanded existing remote sensing image-text pairs to build a multi-round conversation dataset with command-following capabilities[kuckreja2024geochat]. BITA introduced a lightweight Fourier Transformer structure to enhance image-text interaction[10415446], while BAN developed a foundational model for remote sensing change detection tasks, leading to improved task performance[10438490]. Multimodal modeling for SAR images is still in its infancy.
In contrast, multimodal foundation model construction for SAR is still in its infancy: early attempts such as SARLANG[wei2025sarlang] and SARCLIP[jiang2026sarclip], which generates text from detection annotations, and SSL-LIP[10642118], which employs dual-stage self-supervised training with limited labels, have demonstrated feasibility but remain constrained by small-scale, template-based text construction and sparse semantics. Therefore, there is an urgent need to construct a multimodal foundation model adapted to the imaging characteristics of SAR images.
@@ 3.1 FUSAR-GEOVL-1M Image Data
As shown in Fig. 4, the FUSAR-GEOVL-1M dataset presents a clear distribution of data, sample proportions, and representative examples. It is noteworthy that this dataset is built upon several task-oriented SAR datasets previously proposed by our team—namely, FUSAR-MAP[shi2021object], FUSAR-SHIP[hou2020fusar], and FUSAR-AIR[Qian2024Thesis]—which have undergone large-scale expansion and data restructuring. These enhancements have significantly improved the dataset’s coverage, semantic density, and adaptability to downstream tasks. The aforementioned datasets have been downloaded over 6,000 times and are widely cited and applied in the field of SAR interpretation. FUSAR-GEOVL-1M represents the systematic accumulation of our team’s long-term research in SAR image understanding and provides a strong data foundation for multimodal modeling in the SAR domain.
@@ 3.3 FUSAR-KLIP Overall Framework
Specifically, FUSAR-KLIP adopts a standardized dual-encoder architecture to separately model visual and textual representations. The visual encoder employs a vision transformer, where SAR images are divided into fixed-size patches and processed through linear projection and positional encoding before being passed to the Transformer. This yields a uniform visual embedding vector: $f_{v}\in\mathbb{R}^{d}$. The text encoder is based on BERT, which tokenizes and embeds the input text, producing a corresponding embedding vector: $f_{t}\in\mathbb{R}^{d}$ through a multi-layer Transformer. These embeddings are aligned within a shared semantic space. ViT and BERT, both Transformer-based, are widely adopted in multimodal frameworks (e.g., BLIP, ALBEF, OFA[li2021align],[Liu2025Causal]) due to their strong alignment capability, generalization, and downstream task compatibility, making them well-suited as foundational encoders for our model.
@@ 4.1 Implementation Details
ChangeClip[dong2024changeclip] ViT-Base ISPRS-2024 59,246
@@
Geochat[kuckreja2024geochat] ViT-Large CVPR-2024 141,246
VHM[pang2025vhm] ViT-Large AAAI-2025 1,390,405
@@
CLIP*[radford2021learning] ViT-Base ICML-2021 400 M
BLIP*[li2022blip] ViT-Base ICML-2022 100 M
@@ 4.2 Analysis of FUSAR-GEOVL-1M Text
To further compare semantic expressiveness, we selected a subset of FUSAR-GEOVL-1M with manual target annotations and generated templated descriptions using the method in RemoteCLIP [10504785]. As shown in Fig. 10, t-SNE visualization reveals that FUSAR-GEOVL-1M texts exhibit tighter clustering aligned with image features, whereas templated texts are more dispersed, indicating semantic sparsity. Additionally, as illustrated in Fig. 11, we compared and analyzed multiple remote sensing multimodal datasets using two metrics: the text semantic richness index (MTLD[mccarthy2005assessment]) and token count. Overall, FUSAR-GEOVL-1M significantly outperforms existing mainstream datasets in both semantic richness and token count.

### v4

@@ 1 Introduction
In recent years, foundational models built around the Transformer have established a dominant position in the field of computer vision, demonstrating outstanding general representation capabilities through self-supervised paradigms such as masked image modeling or contrastive learning[15]. Particularly, the advent of vision-language models have pioneered a new paradigm of cross-modal semantic alignment, enabling models to transcend simple visual feature extraction and achieve high-level cognitive understanding[65]. Inspired by this, a number of foundational models targeting the characteristics of Earth observation have rapidly emerged in the field of remote sensing. To address the need for high-precision sensing, RingMo[10] and SpectralGPT[15] have achieved deep adaptation to small targets and multispectral features through customized masked modeling and 3D spectral reconstruction. Meanwhile, in order to bridge the gap between low-level pixels and high-level semantics, GeoRSClip[75] and RemoteCLIP[37] introduce natural language alignment to enhance the inference depth of the model. To address the challenges of wide-area generation and data sparsity, MetaEarth[72] and AlphaEarth[4] have introduced generative diffusion and spatiotemporal embedding field techniques, successfully extending the model’s capabilities from passive interpretation to active simulation and multi-source assimilation[22],[21], forming a diversified development pattern centered around the core needs of Earth observation[19],[80],[33],[54],[39].
While the aforementioned work has achieved record-breaking performance metrics on specific tasks, its theoretical foundation still faces significant challenges. As Muttenthaler et al. pointed out, a fundamental misalignment exists between intelligent model representations and human cognition: existing deep neural networks, although excelling in local feature matching, often fail to capture hierarchical knowledge structures like humans[45]. This means that current models lack the deep cognitive abilities that align with human logic, resulting in significantly limited robustness and generalization when faced with out-of-distribution data[42].
@@
This cognitive disparity has further evolved into a dual barrier of modal heterogeneity and mechanism intricacy in the field of synthetic aperture radar (SAR)[20]. Although SAR possesses irreplaceable all-weather strategic value, its imaging, based on microwave coherent mechanisms, is heavily affected by speckle noise and geometric distortion[67]. When existing foundation models are migrated to the SAR domain, they struggle to understand the geometric structure and physical properties of ground objects through non-intuitive scattering phenomena, leading to a sharp decline in generalization performance[79]. Therefore, leveraging the advantages of cross-modal artificial intelligence to enhance cognitive capabilities through vision-language models and constructing a multimodal foundation model that can adapt to SAR imaging characteristics and possess geospatial cognitive capabilities has become a key scientific task in the field of remote sensing intelligence.
@@
1) Geospatial Priors Remain Underutilized: Geographic information, a core element in Earth sciences, is crucial for remote sensing. However, most current SAR interpretation studies remain at the visual level, neglecting geographic attributes. This stems from the extension of natural image paradigms—focused mainly on texture and structure[40]—as well as the closed nature of SAR data sources, with many studies still relying on outdated datasets like MSTAR from the 1990s[1]. As shown in Table I, our review of 16 mainstream SAR datasets across tasks such as detection, segmentation, and recognition reveals a widespread omission of geographic metadata. This absence hinders regional analysis, terrain understanding, and spatial reasoning, ultimately limiting progress toward cognitive-level SAR interpretation.
2) Knowledge Bottleneck: Existing remote sensing VLM studies face a significant knowledge bottleneck in text construction. On the one hand, approaches that convert detection, recognition, or segmentation annotations into templated text [37],[64] generate descriptions that are semantically sparse and structurally repetitive, covering only shallow attributes (e.g., category and location) while neglecting complex elements such as landform, regional function, or spatial structure. This low-information text causes different images to collapse into similar semantic spaces, weakening alignment quality and fine-grained discrimination, as shown in Fig. 2(a). On the other hand, automatic text acquisition strategies, which are effective in natural image or optical remote sensing domains (e.g., web crawling, video subtitles, or AI-generated captions), fail in the SAR domain due to speckle noise and the need for expert knowledge. The resulting AI-generated text is often semantically vague or distorted Fig. 2(b). Together, these issues highlight that SAR multimodal foundational model suffers from a dual bottleneck: text derived from manual annotations lacks semantic depth, and reliable large-scale automatic acquisition remains infeasible, severely restricting scalability in open-world applications.
@@
FUSAR-MAP[52] 2021 45 C Single 1 1,024 610 TIF Segmentation
SSDD[68] 2021 633 C,X Multi $\sim1$5 500 1,160 JPG Detection
@@
FUSAR-SHIP[16] 2022 271 C Single 1 512 16,144 TIFF Detection
@@
SAMPLE[25] 2019 14 X Single 0.3 128 2,732 PNG Classification
@@
PolarSAR-Seg[63] 2022 36 C Multi 8 512 500 TIFF Segmentation
@@
In terms of text data acquisition, this paper introduces a hierarchical cognitive chain-of-thought (HCoT) instruction to simulate the human interpretation process of SAR imagery. HCoT guides GPT-4.1 [46] to progressively incorporate multi-dimensional knowledge—such as geographic context, regional priors, SAR imaging principles, and target scale perception—enabling semantic reasoning from global to local, and from general to domain-specific knowledge. Furthermore, the constructed multi-scale image semantics driving mechanism guides the model to generate semantic expressions that connect the upper and lower scales on large, medium, and small scale images, and establish cognitive coherence. Under HCoT guidance, the language model significantly enhances its SAR interpretation ability, generating more informative and coherent textual descriptions to support multimodal representation learning.
In terms of multimodal alignment modeling, this paper proposes FUSAR-KLIP, a dual-tower multimodal framework employing Vision Transformer (ViT)[12] and BERT[9] as visual and language encoders. The model jointly optimizes image-text contrastive loss (ITC), image-text matching loss (ITM), and masked language modeling loss (MLM) to construct a unified cross-modal embedding space bridging low-level perception and high-level semantics. To mitigate semantic deviations in HCoT-guided generated text, we introduce a self-consistent iterative optimization (SCIO) module that enhances alignment accuracy and stability through a closed-loop self-supervised strategy of screening, filtering, and reconstruction.
@@ 2.1 Paradigm Evolution of Remote Sensing Interpretation
In the supervised learning stage, large-scale manual annotations enabled significant progress in recognition, detection, and segmentation tasks, but the prohibitive cost and limited semantic coverage of annotations severely constrained generalization across diverse scenes[52].
Unimodal self-supervised learning (SSL) alleviated this bottleneck by mining latent structures from unlabeled data, producing robust representations with minimal fine-tuning and laying the foundation for domain-specific models[77],[35]. The RingMo series pioneered the introduction of mask learning into remote sensing, optimizing feature extraction of dense small targets in remote sensing images through targeted masking strategies[10]. RingMo-Aerial further introduced a frequency-enhanced attention mechanism and affine transformation contrastive learning, solving the challenges of tilted viewpoints and multi-scale occlusion in aerial remote sensing. Addressing the 3D characteristics of multispectral data, SpectralGPT innovatively employs 3D masking and multi-target reconstruction strategies, overcoming the limitation of processing only RGB data and achieving effective capture of spectral sequence information[15]. In the generative direction, MetaEarth, through a self-cascaded framework and noise sampling strategy, achieved global-scale, unbounded resolution remote sensing image generation, driving the leap from perception to simulation[72]. RS-vHeat takes a different approach, using a heat conduction physics model to simulate feature diffusion, significantly improving computational efficiency while maintaining the global receptive field[17]. Furthermore, AlphaEarth proposed the concept of Embedding Fields, addressing the label sparsity problem by assimilating multi-source spatiotemporal data[4]. Nevertheless, unimodal SSL remains largely restricted to low-level perceptual cues such as texture and geometry, offering limited capacity for higher-order semantic reasoning[38],[35].
@@ 2.2 Research on Unimodal Self-supervised Foundational Models in SAR Images
SARATR-X uses a two-step SSL method with multi-scale gradient features to establish a high-performance SAR image target recognition foundational model[29]. Yang et al. proposed the SARDet-CL, combining feature enhancement with a SSL method constrained by imaging mechanisms, and achieved advanced performance in downstream detection tasks[71]. Li et al. proposed SAR-JEPA, which constructed a self-supervised pre-training foundational model for SAR target recognition tasks from the perspective of overcoming speckle noise[30]. Ren et al. introduced multi-image factor SSL to promote directional feature learning and obtain generalized features, enhancing the performance in terrain classification tasks[51]. Wang et al. constructed a SAR multi-task foundation model based on cross-domain continuous pre-training[60]. Pei et al. proposed a two-stage SAR image pre-training method based on SSL to improve the accuracy of target classification[48]. MSFA proposed a multi-stage filter augmentation pre-training framework for large-scale RGB and SAR data, which performed well when transferred to detection tasks[34].
@@ 2.3 Research on Multimodal VLMs in Remote Sensing Field
Existing work primarily focuses on optical imagery[58]. For instance, RemoteCLIP is based on the CLIP framework by converting class labels in public remote sensing datasets into templated text, establishing a foundational model that achieves leading performance across many typical tasks[37]. SkyEyeGPT developed a multimodal remote sensing command dataset and designed a two-stage tuning strategy to enhance conversational capabilities[73]. VHM generated the HqDC dataset based on the large language model Gemini, easing the issue of model hallucination[47]. SkyScript utilized structured geographic information from OpenStreetMap to generate semantic text for optical images, improving the quality of cross-modal alignment[62]. GeoChat expanded existing remote sensing image-text pairs to build a multi-round conversation dataset with command-following capabilities[24]. BITA introduced a lightweight Fourier Transformer structure to enhance image-text interaction[69], while BAN developed a foundational model for remote sensing change detection tasks, leading to improved task performance[28]. Multimodal modeling for SAR images is still in its infancy.
In contrast, multimodal foundation model construction for SAR is still in its infancy: early attempts such as SARLANG[64] and SARCLIP[23], which generates text from detection annotations, and SSL-LIP[70], which employs dual-stage self-supervised training with limited labels, have demonstrated feasibility but remain constrained by small-scale, template-based text construction and sparse semantics. Therefore, there is an urgent need to construct a multimodal foundation model adapted to the imaging characteristics of SAR images.
@@ 3.1 FUSAR-GEOVL-1M Image Data
As shown in Fig. 4, the FUSAR-GEOVL-1M dataset presents a clear distribution of data, sample proportions, and representative examples. It is noteworthy that this dataset is built upon several task-oriented SAR datasets previously proposed by our team—namely, FUSAR-MAP[52], FUSAR-SHIP[16], and FUSAR-AIR[49]—which have undergone large-scale expansion and data restructuring. These enhancements have significantly improved the dataset’s coverage, semantic density, and adaptability to downstream tasks. The aforementioned datasets have been downloaded over 6,000 times and are widely cited and applied in the field of SAR interpretation. FUSAR-GEOVL-1M represents the systematic accumulation of our team’s long-term research in SAR image understanding and provides a strong data foundation for multimodal modeling in the SAR domain.
@@ 3.3 FUSAR-KLIP Overall Framework
Specifically, FUSAR-KLIP adopts a standardized dual-encoder architecture to separately model visual and textual representations. The visual encoder employs a vision transformer, where SAR images are divided into fixed-size patches and processed through linear projection and positional encoding before being passed to the Transformer. This yields a uniform visual embedding vector: $f_{v}\in\mathbb{R}^{d}$. The text encoder is based on BERT, which tokenizes and embeds the input text, producing a corresponding embedding vector: $f_{t}\in\mathbb{R}^{d}$ through a multi-layer Transformer. These embeddings are aligned within a shared semantic space. ViT and BERT, both Transformer-based, are widely adopted in multimodal frameworks (e.g., BLIP, ALBEF, OFA[27],[41]) due to their strong alignment capability, generalization, and downstream task compatibility, making them well-suited as foundational encoders for our model.
@@ 4.1 Implementation Details
ChangeClip[11] ViT-Base ISPRS-2024 59,246
@@
Geochat[24] ViT-Large CVPR-2024 141,246
VHM[47] ViT-Large AAAI-2025 1,390,405
@@
CLIP*[50] ViT-Base ICML-2021 400 M
BLIP*[26] ViT-Base ICML-2022 100 M
@@ 4.2 Analysis of FUSAR-GEOVL-1M Text
To further compare semantic expressiveness, we selected a subset of FUSAR-GEOVL-1M with manual target annotations and generated templated descriptions using the method in RemoteCLIP [37]. As shown in Fig. 10, t-SNE visualization reveals that FUSAR-GEOVL-1M texts exhibit tighter clustering aligned with image features, whereas templated texts are more dispersed, indicating semantic sparsity. Additionally, as illustrated in Fig. 11, we compared and analyzed multiple remote sensing multimodal datasets using two metrics: the text semantic richness index (MTLD[44]) and token count. Overall, FUSAR-GEOVL-1M significantly outperforms existing mainstream datasets in both semantic richness and token count.